# CIVIL: Causal and Intuitive Visual Imitation Learning

### v2

@@ V Simulations
Figure 6: Results for the real-world experiments in Section VI. We compare our proposed approach (CIVIL) to object-oriented (Task-VIOLA) and language-conditioned (FiLM) approaches in four manipulation tasks: (1) cooking vegetables or meat, (2) pressing a red button, (3) picking a cup, and (4) pulling a bowl to the center of the table. The top row shows examples of training scenarios in each task. We highlight the task-relevant objects in green and one of the distracting objects with a red boundary. During training, the position of the distracting object can be correlated with the relevant object, but this correlation is not present during testing. Example test scenarios are shown in the second row, where the target object appears in an unseen position (except in the cooking task, where the target is fixed). The bottom row shows the success rates of the robot arm. We use a darker shade to denote success in seen scenarios and a lighter shade for unseen scenarios. The performance for all approaches drops as the tasks become more complex. In our experiments, cooking was the easiest task as the pan was fixed, while pulling was the most challenging because it involved two target-oriented subtasks, reaching the bowl and bringing it to the center. CIVIL achieves a significantly higher success rate than the object-oriented and language-conditioned approaches across all real-world tasks.
VI Real-World Experiments
@@
In this section we compare our approach to two visual imitation learning baselines: i) Task-VIOLA: the method from our simulations that receives segmented images of the task-relevant objects, and ii) FiLM [59]: a vision-language baseline that uses language prompts to condition its visual features with an affine transformation. We applied this approach instead of CLIP because we found that the features obtained from a pre-trained CLIP encoder did not work well in our real-world tasks during initial testing. Unlike CLIP, FiLM requires language prompts during both training and testing.
Experimental Setup. We evaluate these methods on a 7-DOF Franka Emika Panda robot arm mounted on a table. The robot uses two Logitech C920 webcams to observe the environment: one serves as a static camera that captures the entire scene, and the other functions as an egocentric camera attached to the end effector of the robot arm. We also use a microphone to record verbal instructions during demonstrations. To indicate relevant poses, we use $3$D-printed cubes with a width of $20$ millimeters as the physical markers. Five faces of the cube have ArUco tags that uniquely identify that marker, while the sixth face has a reusable adhesive. If the robot detects more than one face of a marker, we take the average of their positions.
@@
Demonstrations The training demonstrations are provided by an expert human using a Logitech joystick. Before performing the task, the expert attaches markers to the target objects. During each demonstration, the robot records the static and egocentric RGB images $y_{env}$, $y_{ego}$ which are resized to $200\times 200$, the $8$-dimensional robot state $x$ which includes $7$ joint angles and one binary gripper state, and a $8$-dimensional action $u$. The action is a $7$-dimensional joint velocity and a binary gripper action. The robot also tracks the positions $b$ of the markers with the static camera. However, the markers are not detected in every frame so we only obtain marker poses for a subset of the images collected during demonstrations. The demonstrator provides verbal instructions $l$ (e.g., scoop the meat or stir the vegetables) which are recorded with a microphone and then transcribed to text by a speech recognition model [60]. We use an open-world video segmentation model DEVA [51] to obtain bounding boxes ${}\{\mathcal{B}\}$ from text prompts. Lastly, we in-paint the markers from the images using OpenCV’s inpainting tools.
Results. Our results are summarized in Figure 6. We trained the methods with $20$ expert demonstrations in each task except button pushing, for which we provided $10$ demonstrations. We then tested the methods in several scenarios that reasonably covered all distinct object configurations in each task. Specifically, for tasks numbered $1$ to $4$, we had $40$, $9$, $16$, and $24$ test scenarios, respectively. We measured success based on whether the robot completed the intended task correctly. For instance, if the robot gripper touched anywhere on the button in the pressing task, it was recorded as a success. But if the robot missed the button, it was a failure. The success rates are averaged over $3$ training and evaluation runs.
@@
VII User Study
Now that we have evaluated our algorithm in simulated and real-world tasks with examples provided by a human expert, we will assess whether it is also easy for everyday humans to convey their reasoning while demonstrating the task. Specifically, we conduct a study to determine if users can intuitively place markers and seamlessly issue language prompts without impacting the quality of their demonstrations. We acknowledge that deploying these inputs requires additional time, which may reduce the time left for users to provide examples. Hence, we also evaluate whether our high-level insight of communicating the key features (why) along with the demonstrations (what) is practically advantageous when users have a fixed amount of time to teach the robot. We compare our approach to the behavior cloning (BC) baseline introduced in simulations. The difference between these methods captures our proposed re-framing of imitation learning: BC learns only from what the human does, while CIVIL enables the human to also convey why they are showing those actions.
Figure 7: User study results from Section VII. (Left) Subjective ratings on a $7$-point Likert scale, where higher values indicate agreement with the statements in Table I. Users subjectively perceived the process of placing markers and explaining actions to be intuitive. Users also reported that they were able to seamlessly include these additional steps into their teaching process. (Middle) Objectively, augmenting the demonstrations with feature supervision (CIVIL) led to a significantly better performance than simply providing more action demonstrations to the robot (BC). Here we use a darker shade to denote success in releasing the cup without accidentally toppling it over. (Right) We show an example rollout of both approaches. In this example CIVIL accurately takes the cup to the coffee machine after picking it up, while BC mistakenly takes the cup to the wrong location.
@@ VII User Study
Participants and Procedure. We recruited $10$ participants ($1$ female, average age $24.4{\pm}3.7$) from the university’s student population. Participants received monetary compensation for their time and provided written consent according to university guidelines (IRB #23-1237).
At the beginning of the study we showed participants a video of an expert demonstration and gave them $5$ minutes to practice teleoperating the robot using the joystick. During this practice session we also instructed participants on how to attach markers and give language prompts. In particular, we told users that markers should be attached to objects of interest such that they are visible to the robot’s camera. Our study followed a within-subjects design where participants provided data in two rounds, once with markers and language, and once without the additional inputs. In each round, users had $5$ total minutes to provide as many demonstrations as possible. This included the time required to place markers and collect any play data. After the trials participants answered a survey (see Table I) to rate their teaching experiences. We counterbalanced the order so that half of the participants worked with CIVIL first, and the other half started with BC.
TABLE I: Survey with two 7-point Likert scales for assessing the intuitiveness of using markers and language, and the ease of incorporating these inputs alongside task demonstrations.
@@
Dependent Variables. To assess the ease of deploying our approach we consider two subjective attributes: Intuitive and Seamless. We measure these attributes through the 7-point Likert scale survey shown in Table I. Users respond to each item in this survey with an agreement rating from $1$ to $7$, where $1$ is strongly disagree and $7$ is strongly agree. Higher ratings indicate participants found it intuitive to use markers and language, and they could seamlessly integrate these inputs into their demonstrations. We evaluate the robot’s objective performance through the success rate of the learned policy.
@@
H1. Users will find teaching robots with CIVIL (i.e., showing demonstrations with markers and language) to be just as intuitive and seamless as providing action demonstrations for standard BC.
@@
Training and Testing. In a time window of $5$ minutes users provided an average of ${\sim}11$ demonstrations without any additional inputs, and ${\sim}9$ demonstrations when also working with markers and language. We aggregated the data provided by users into two datasets: i) $\mathcal{D}_{BC}$ which includes the states $x$, images $y$, and actions $u$ from the baseline round, and ii) $\mathcal{D}_{CIVIL}$ which includes marker readings $b$ and language prompts $l$ along with the $(x,y,u)$ samples from our proposed round. We also processed the user’s language commands to extract the relevant objects. Specifically, we computed the cosine similarity between the text transcribed by Whisper [60] and a pre-defined library containing descriptions for all objects in the environment. For example, “coffee machine” and “black Keurig” were mapped to “black coffee maker”.
@@
Results. Figure 7 summarizes our user study outcomes.
Overall, users reported that they found it intuitive to deploy markers and speak language commands while teleoperating the robot. To evaluate their subjective responses, we combined ratings for the survey questions into two scores: one for intuitive and one for seamless. T-tests indicated that the average user scores for the Intuitive ($t(9)=12.99$, $p<0.001$) and Seamless ($t(9)=4.34$, $p<0.001$) scales were significantly higher than the neutral score of $4$. We note that we did not physically show users how to attach markers; we only gave them verbal instructions during the practice round. Therefore, this result indicates that it was easy for users to understand, remember, and implement our data collection procedure. It also supports our hypotheses H1. However, we caveat this result with the awareness that $8$ of our $10$ participants stated they had previously interacted with robots, which may have helped them comprehend how robots learn from visual observations.
@@
Lastly, when taking a closer look at where our approach was superior to standard behavior cloning, we found that CIVIL was significantly more successful in picking up and releasing the cup than BC. For instance, the robot failed to pick up the cup in only $15\%$ of the test scenarios when using CIVIL, compared to $48\%$ when trained with BC. Also, when the robot did manage to pick the cup and take it to the coffee machine, CIVIL was $30\%$ more successful than BC in releasing the cup and moving out without knocking it over. Both these instances represent key states in the task where the robot needs to be the most accurate. This is where training with marker readings helps CIVIL to be more precise than conventional approaches that rely on the robot to extract such positional features without any human guidance.
@@
VIII Conclusion
In this paper we tackle the problem of causal confusion in visual imitation learning by proposing a fundamental shift in the way humans provide demonstrations, and then leveraging that augmented data to explain the human’s actions. We prove that it is difficult for robots to learn from visual demonstrations because of the high-dimensionality of the inputs and the spurious correlations in their datasets. To address this challenge, we propose that humans supplement their action demonstrations with additional cues that reveal their decision-making process. Specifically, we enable humans to deploy physical markers and utter natural language instructions to intuitively convey task-relevant positions and objects that inform their decisions.
@@ VIII Conclusion
Limitations and Future Work. This work is a step towards maximizing what robots can learn from human examples. Our current approach has some limitations. For instance, we rely on humans to mark or mention the relevant objects. This may lead to errors when teaching tasks that contain several relevant components — humans could forget an essential object or mistakenly mention an irrelevant object. Future work should account for such potential human errors to prevent the robot from learning incomplete or non-causal representations. A possible way to mitigate this issue would be to actively remind and interact with users throughout the demonstration process. Another limitation of our work is that we only use verbal commands to identify the task-relevant objects. However, human instructions often contain additional insights such as qualitative descriptions of the demonstrated action (e.g., “go straight to” or “place carefully under”). Leveraging these latent signals can help robots make full use of the human’s inputs and further accelerate the learning process.

### v3

@@ 5 Simulations
Figure 7: Results for the real-world experiments in Section 6. We compare our proposed approach (CIVIL) to object-oriented (Task-VIOLA) and language-conditioned (FiLM) approaches in four manipulation tasks: (1) cooking vegetables or meat, (2) pressing a red button, (3) picking a cup, and (4) pulling a bowl to the center of the table. The top row shows examples of training scenarios in each task. We highlight the task-relevant objects in green and one of the distracting objects with a red boundary. During training, the position of the distracting object can be correlated with the relevant object, but this correlation is not present during testing. Example test scenarios are shown in the second row, where the target object appears in an unseen position (except in the cooking task, where the target is fixed). The bottom row shows the success rates of the robot arm. We use a darker shade to denote success in seen scenarios and a lighter shade for unseen scenarios. The performance for all approaches drops as the tasks become more complex. In our experiments, cooking was the easiest task as the pan was fixed, while pulling was the most challenging because it involved two target-oriented subtasks, reaching the bowl and bringing it to the center. CIVIL achieves a significantly higher success rate than the object-oriented and language-conditioned approaches across all real-world tasks.
6 Real-World Experiments
@@
In this section we compare our approach to two visual imitation learning baselines: i) Task-VIOLA: the method from our simulations that receives segmented images of the task-relevant objects, and ii) FiLM perez2018film : a vision-language baseline that uses language prompts to condition its visual features with an affine transformation. We applied this approach instead of CLIP because we found that the features obtained from a pre-trained CLIP encoder did not work well in our real-world tasks during initial testing. Unlike CLIP, FiLM requires language prompts during both training and testing.
Experimental Setup. We evaluate these methods on a 7-DOF Franka Emika Panda robot arm mounted on a table. The robot uses two Logitech C920 webcams to observe the environment: one serves as a static camera that captures the entire scene, and the other functions as an egocentric camera attached to the end effector of the robot arm. We also use a microphone to record verbal instructions during demonstrations. To indicate relevant poses, we use 33D-printed cubes with a width of $20$ millimeters as the physical markers. Five faces of the cube have ArUco tags that uniquely identify that marker, while the sixth face has a reusable adhesive. If the robot detects more than one face of a marker, we take the average of their positions.
@@
Demonstrations The training demonstrations are provided by an expert human using a Logitech joystick. Before performing the task, the expert attaches markers to the target objects. During each demonstration, the robot records the static and egocentric RGB images $y_{env}$, $y_{ego}$ which are resized to $200\times 200$, the $8$-dimensional robot state $x$ which includes $7$ joint angles and one binary gripper state, and a $8$-dimensional action $u$. The action is a $7$-dimensional joint velocity and a binary gripper action. The robot also tracks the positions $b$ of the markers with the static camera. However, the markers are not detected in every frame so we only obtain marker poses for a subset of the images collected during demonstrations. The demonstrator provides verbal instructions $l$ (e.g., scoop the meat or stir the vegetables) which are recorded with a microphone and then transcribed to text by a speech recognition model radford2023robust . We use an open-world video segmentation model DEVA cheng2023tracking to obtain bounding boxes $\{\mathcal{B}\}$ from text prompts. Lastly, we in-paint the markers from the images using OpenCV’s inpainting tools.
Results. Our results are summarized in Figure 7. We trained the methods with $20$ expert demonstrations in each task except button pushing, for which we provided $10$ demonstrations. We then tested the methods in several scenarios that reasonably covered all distinct object configurations in each task. Specifically, for tasks numbered $1$ to $4$, we had $40$, 99, $16$, and $24$ test scenarios, respectively. We measured success based on whether the robot completed the intended task correctly. For instance, if the robot gripper touched anywhere on the button in the pressing task, it was recorded as a success. But if the robot missed the button, it was a failure. The success rates are averaged over $3$ training and evaluation runs.
@@
7 User Study
Now that we have evaluated our algorithm in simulated and real-world tasks with examples provided by a human expert, we will assess whether it is also easy for everyday humans to convey their reasoning while demonstrating the task. Specifically, we conduct a study to determine if users can intuitively place markers and seamlessly issue language prompts without impacting the quality of their demonstrations. We acknowledge that deploying these inputs requires additional time, which may reduce the time left for users to provide examples. Hence, we also evaluate whether our high-level insight of communicating the key features (why) along with the demonstrations (what) is practically advantageous when users have a fixed amount of time to teach the robot. We compare our approach to the behavior cloning (BC) baseline introduced in simulations. The difference between these methods captures our proposed re-framing of imitation learning: BC learns only from what the human does, while CIVIL enables the human to also convey why they are showing those actions (i.e., the human conveys which robot and environment features their policy depends upon).
Figure 8: User study results from Section 7. (Left) Subjective ratings on a 77-point Likert scale, where higher values indicate agreement with the statements in Table 1. Users subjectively perceived the process of placing markers and explaining actions to be intuitive. Users also reported that they were able to seamlessly include these additional steps into their teaching process. (Middle) Objectively, augmenting the demonstrations with feature supervision (CIVIL) led to a significantly better performance than simply providing more action demonstrations to the robot (BC). Here we use a darker shade to denote success in releasing the cup without accidentally toppling it over. (Right) We show an example rollout of both approaches. In this example CIVIL accurately takes the cup to the coffee machine after picking it up, while BC mistakenly takes the cup to the wrong location.
@@ 7 User Study
Participants and Procedure. We recruited $10$ participants ($1$ female, average age $24.4{\pm}3.7$) from Virginia Tech’s student population. Participants received monetary compensation for their time and provided informed written consent according to university guidelines (IRB #23-1237).
At the beginning of the study we showed participants a video of an expert demonstration and gave them $5$ minutes to practice teleoperating the robot using the joystick. During this practice session we also instructed participants on how to attach markers and give language prompts. In particular, we told users that markers should be attached to objects of interest such that they are visible to the robot’s camera. Our study followed a within-subjects design where participants provided data in two rounds, once with markers and language, and once without the additional inputs. In each round, users had $5$ total minutes to provide as many demonstrations as possible. This included the time required to place markers and collect any play data. It is important to note that the markers only need to be attached once at the beginning of one round, and they can remain in place throughout all subsequent demonstrations. Therefore, attaching the markers does not add significantly to the total time required or the human effort involved. After the trials participants answered a survey (see Table 1) to rate their teaching experiences. We counterbalanced the order so that half of the participants worked with CIVIL first, and the other half started with BC.
Table 1: Survey with two 7-point Likert scales for assessing the intuitiveness of using markers and language, and the ease of incorporating these inputs alongside task demonstrations.
@@
Dependent Variables. To assess the ease of deploying our approach we consider two subjective attributes: Intuitive and Seamless. We measure these attributes through the 7-point Likert scale survey shown in Table 1. Users respond to each item in this survey with an agreement rating from $1$ to $7$, where $1$ is strongly disagree and $7$ is strongly agree. Higher ratings indicate participants found it intuitive to use markers and language, and they could seamlessly integrate these inputs into their demonstrations. We evaluate the robot’s objective performance through the success rate of the learned policy.
@@
H1. Users will find teaching robots with CIVIL (i.e., showing demonstrations with markers and language) to be just as intuitive and seamless as providing demonstrations for standard BC.
@@
Training and Testing. In a time window of $5$ minutes users provided an average of ${\sim}11$ demonstrations without any additional inputs, and ${\sim}9$ demonstrations when also working with markers and language. We aggregated the data provided by users into two datasets: i) $\mathcal{D}_{BC}$ which includes the states $x$, images $y$, and actions $u$ from the baseline round, and ii) $\mathcal{D}_{CIVIL}$ which includes marker readings $b$ and language prompts $l$ along with the $(x,y,u)$ samples from our proposed round. We also processed the user’s language commands to extract the relevant objects. Specifically, we computed the cosine similarity between the text transcribed by Whisper radford2023robust and a pre-defined library containing descriptions for all objects in the environment. For example, “coffee machine” and “black Keurig” were mapped to “black coffee maker”.
@@
Results. Figure 8 summarizes our study outcomes.
Overall, users reported that they found it intuitive to deploy markers and speak language commands while teleoperating the robot. To evaluate their subjective responses, we combined ratings for the survey questions into two scores: one for intuitive and one for seamless. T-tests indicated that the average user scores for the Intuitive ($t(9)=12.99$, $p<0.001$) and Seamless ($t(9)=4.34$, $p<0.001$) scales were significantly higher than the neutral score of $4$. We note that we did not physically show users how to attach markers; we only gave them verbal instructions during the practice round. Therefore, this result indicates that it was easy for users to understand, remember, and implement our data collection procedure. It also supports our hypotheses H1. However, we caveat this result with the awareness that $8$ of our $10$ participants stated they had previously interacted with robots, which may have helped them comprehend how robots learn from visual observations and provide more informed training data.
@@
Lastly, when taking a closer look at where our approach was superior to standard behavior cloning, we found that CIVIL was significantly more successful in picking up and releasing the cup than BC. For instance, the robot failed to pick up the cup in only $15\%$ of the test scenarios when using CIVIL, compared to $48\%$ when trained with BC. Also, when the robot did manage to pick the cup and take it to the coffee machine, CIVIL was $30\%$ more successful than BC in releasing the cup and moving out without knocking it over. Both these instances represent key states in the task where the robot needs to be the most accurate. This is where training with marker readings helps CIVIL to be more precise than conventional approaches that rely on the robot to extract such positional features without any human guidance. We also hypothesize that the expressiveness of natural language commands helped non-expert users. When the robot purely learns from motion demonstrations, any errors in these movements can lead to confusion. But when the robot reasons over the associated language and markers, CIVIL enables the robot to correctly parse the relevant features, even if the human’s motions are imperfect.
@@
8 Conclusion
In this paper we tackle the problem of causal confusion in visual imitation learning by proposing a fundamental shift in the way humans provide demonstrations, and then leveraging that augmented data to explain the human’s actions. Given just action demonstrations and high-dimensional visual observations, robots can struggle to autonomously extract the correct feature representations. Without these representations, we analytically and experimentally show that robots may learn to condition their policies on extraneous or spuriously-correlated data, leading to out-of-distribution failures. To address this challenge, we propose that humans supplement their action demonstrations with additional cues that reveal their decision-making process. Specifically, we enable humans to deploy physical markers and utter natural language instructions to intuitively convey task-relevant positions and objects that form part of the desired feature representation.
@@ 8 Conclusion
Limitations and Future Work. This work is a step towards maximizing what robots can learn from human examples. However, our current approach has some limitations. For instance, we rely on humans to mark or mention the relevant objects. This may lead to errors when teaching tasks that contain several relevant components — humans could forget an essential object or mistakenly mention an irrelevant object. Future work should account for such potential human errors to prevent the robot from learning incomplete or non-causal representations. A possible way to mitigate this issue would be to actively remind and interact with users throughout the demonstration process. Another limitation of our work is that we only use verbal commands to identify the task-relevant objects. However, human instructions often contain additional insights such as qualitative descriptions of the demonstrated action (e.g., “go straight to” or “place carefully under”). Leveraging these latent signals can help robots make full use of the human’s inputs and further accelerate the learning process.
9 Declarations
Funding. This research was supported in part by the NSF (Grant Number $2337884$).
Conflict of Interest. The authors declare that they have no conflicts of interest.
Ethical Statement. All physical experiments that relied on interactions with humans were conducted under university guidelines and followed the protocol of Virginia Tech IRB #23\#23-$1237$.
Author Contribution. Y.D. and R.S. led the algorithm development for visual feature extraction. Y.D. contributed to the use of physical markers. R.S. led the setup of the simulation environment. H.N. led the development of the theoretical background. H.N. and D.L. wrote the first manuscript draft. Y.D. and R.J. ran the simulations and R.S. conducted the physical experiments. S.S. and C.N. provided valuable assistance and support throughout the project. D.L. supervised the project, helped develop the method, and edited the manuscript.